# FREE-Merging: Fourier Transform for Efficient Model Merging

### v2

@@ E.2 Merging Language Models
Task Arithmetic 67.62 53.38 78.24 66.41
Task Arithmetic (w/ DARE) 67.82 52.66 82.83 67.77
Ties-Merging 68.27 50.01 84.10 67.46
Ties-Merging (w/ DARE) 69.32 53.07 81.19 67.86
Breadcrumbs 68.24 52.88 79.20 66.77
PCB-Merging 67.23 52.48 80.37 66.69
@@
RegMean 67.89 52.45 82.24 67.52
EMR-Merging 67.82 55.01 90.13 70.98
Twin-Merging 68.32 55.76 90.98 71.68
@@
In this section, we present the complete experimental results of merging language models.
First, we show the merging results for RoBERTa as the pre-trained model, fine-tuned separately on the eight tasks of GLUE, as summarized in Table 15. Our method achieves optimal results on nearly all datasets, demonstrating its generalization capability on language models. Additionally, Table 16 presents the merging performance on 11 language classification tasks using T0-3B as the pre-trained model and $IA^{3}$ as the fine-tuning method. Our approach similarly shows an advantage here. Finally, Table 17 displays the merging performance for QWEN-14B as the pre-trained model and LoRA as the fine-tuning method across three generative tasks. Both our FR-Merging and FREE-Merging exhibit significant advantages over competing methods in each respective category.
Overall, our proposed methods demonstrate strong generalization capabilities on language models.

### v3

@@ E.2 Merging Language Models
Task Arithmetic [21] 67.62 53.38 78.24 66.41
Task Arithmetic (w/ DARE) [53] 67.82 52.66 82.83 67.77
Ties-Merging [48] 68.27 50.01 84.10 67.46
Ties-Merging (w/ DARE) [53] 69.32 53.07 81.19 67.86
Breadcrumbs [7] 68.24 52.88 79.20 66.77
PCB-Merging [12] 67.23 52.48 80.37 66.69
@@
RegMean [23] 67.89 52.45 82.24 67.52
EMR-Merging [20] 67.82 55.01 90.13 70.98
Twin-Merging [29] 68.32 55.76 90.98 71.68
@@
In this section, we present the complete and detailed experimental results of merging a diverse set of language models, each fine-tuned under different settings, across multiple widely-used benchmark datasets that cover both classification and generative tasks.
First, we present the merging results using RoBERTa as the pre-trained model, which was fine-tuned individually on the eight tasks included in the GLUE benchmark. These results are summarized in Table 15. Our proposed method achieves the best performance on nearly all of the datasets, highlighting its strong generalization capability when applied to language models. In addition, Table 17 reports the merging results across 11 language classification tasks, where T0-3B serves as the pre-trained model and $IA^{3}$ is used for fine-tuning. Our method also demonstrates a clear advantage in this setting. Finally, Table 18 provides the results for merging with QWEN-14B as the pre-trained model and LoRA as the fine-tuning technique, evaluated on three generative tasks. In all these experiments, both FR-Merging and FREE-Merging deliver consistently superior performance compared to other existing model merging methods, confirming their effectiveness across various fine-tuning strategies and model types.
Overall, our proposed methods demonstrate strong generalization capabilities when applied to language models, consistently delivering robust performance across a wide range of tasks, architectures, and fine-tuning strategies.